OpenStudy (anonymous):

how do I simplify the square root of 8 - the square root of 50 + 7times the square root of 2

OpenStudy (anonymous):

\[\sqrt{8}-\sqrt{50}+7\sqrt{2}\] \[2\sqrt{2}-5\sqrt{2}+7\sqrt{2}\] \[4\sqrt{2}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome :)